# ReMamba: Equip Mamba with Effective Long-Sequence Modeling

### v2

@@ Abstract
Introduction
@@
Transformers (Vaswani et al. 2017), which form the backbone of most LLMs, encounter substantial challenges when dealing with long texts. The quadratic computational demands and the linear memory costs of the attention mechanism become prohibitive as the text length grows. This complexity poses a significant barrier to effectively modeling long texts, which is crucial for the development of LLMs. To address this, Mamba is proposed as a solution (Gu and Dao 2024). Mamba models utilize a recurrent inference mode that ensures linear time complexity and compress information into the fixed state size, resulting in constant memory demands during inference. Furthermore, Mamba models eliminate the need for positional encoding, theoretically allowing them to handle inputs of any length. Mamba performs competitively against transformers on downsteam tasks. Shortly after, Mamba2 was introduced, simplifying the structured $A$ matrix of Mamba to enable faster training and enlarged state size (Dao and Gu 2024).
@@
Figure 1: A comparison of pretrained Mamba models and Transformers of equivalent size across speed, short-context, and long-context performance metrics. Speed is measured under conditions of 6k input tokens and 1k output tokens. “short scores” represents the average accuracy across six tasks (HellaSwag, PIQA, Arc-E, Arc-C, WinoGrande, OpenbookQA) evaluated within the LM evaluation harness (Gao et al. 2023). “long scores” corresponds to the average scores on the LongBench-E benchmark (Bai et al. 2024). Notably, all LongBench evaluations employ a maximum token length of 2k to align with the model’s training configuration.
@@
Despite these advantages, some studies reveal that Mamba models do not perform as well as expected when dealing with long texts reaching 2k tokens or more (Waleffe et al. 2024). As depicted in Figure 1, our experimental findings reveal that the pretrained Mamba model surpasses pretrained Transformers of comparable size, such as llama2-3b (Geng and Liu 2023), on short-context tasks. Conversely, a substantial performance degradation is observed for Mamba on long-context tasks relative to Transformers. This performance disparity underscores a significant limitation of Mamba models in practical long-context applications.
@@
This long-context deficency issue of Mamba is usually attributed to its RNN-like nature. This kind of architecture exhibits limitations in preserving crucial information from earlier input sequences as the context length increases due to the fixed-size memory (Wen, Dang, and Lyu 2024; Yang et al. 2024b). Hybrid architectures (Lieber et al. 2024; Ren et al. 2024; Park et al. 2024) have sought to mitigate this issue by integrating attention mechanisms from transformers. However, these approaches often lead to decreased computational efficiency and increased memory consumption. A parallel study, DeciMamba (Ben-Kish et al. 2024), also attributes Mamba’s limitations to a restricted effective receptive field and proposes a method for discarding less important tokens in specific layers to extend the length capabilities of Mamba. However, their focus is on improving the length extrapolation ability without training, and they still achieve limited performance.
@@
To improve the long-context performance of Mamba, we introduce ReMamba. The core intention in ReMamba is straightforward. The distant information within Mamba undergoes excessive degradation. An effective compression strategy to condense information and reduce the distance can be of help. Our approach achieves this compression by selecting the top-k hidden states during the first forward pass and leverages Mamba’s selective mechanism to incorporate them into the state space during the second forward pass. ReMamba incurs minimal additional computational overhead (a single extra forward pass). Experimental results demonstrate that our approach significantly improves Mamba’s long-context performance, bringing it close to the performance of transformers. Our ReMamba model achieves a 3.2 improvement over the baseline on LongBench (Bai et al. 2024) and 1.6 improvement on L-Eval (An et al. 2023). Furthermore, our methodology exhibits transferability to Mamba2, yielding a 1.6 improvement on LongBench, offering a broader impact on the Mamba model family.
Preliminaries
This section provides an overview of the Mamba architecture’s development.
State Space Models
State space sequence models (SSMs) are a class of models that take inspiration from classical continuous dynamical systems. To efficiently compute the state transformations, structured SSMs impose specific constraints on the state transition matrix $\hat{A}$. Mamba models are a prime example, employing a diagonal structure for this matrix. In the one-dimensional situation, discrete structured SSMs transform sequences as follows:
$h_{t+1}$ $=\hat{A}h_{t}+\hat{B}x_{t}$ (1)
$y_{t+1}$ $=Ch_{t+1}$
Here $h_{t}\in\mathbb{R}^{N\times 1}$ are state vectors, $\hat{A}\in\mathbb{R}^{N\times N}$ are state transition matrices, $\hat{B}\in\mathbb{R}^{N}$ are input coefficient matrices, $x_{t}\in\mathbb{R}$ are inputs, $C\in\mathbb{R}^{1\times N}$ is an output matrix, and $N$ is the state dimension. For simplicity, matrix multiplications are omitted.
Mamba
@@
The state space model above chooses the time-invariant $\hat{A}$ (state transition matrix) and $\hat{B}$ (input coefficient matrix) thus lacking expressiveness and flexibility. Mamba (Gu and Dao 2024) proposes to make $\hat{A}$ and $\hat{B}$ dynamically depend on inputs.
@@
Furthermore, $\Delta_{t-1}^{l}\in\mathbb{R}^{H^{\prime}}$ is the discrete time step corresponding to the selective mechanism in Mamba, where $H^{\prime}$ is the intermediate hidden size. The continuous and discrete state transformation matrices at layer $l$ are given by $A^{l},\hat{A}^{l}\in\mathbb{R}^{H^{\prime}\times N}$, respectively. The continuous and discrete input coefficient matrices are denoted by $B_{t-1}^{l},\hat{B}_{t-1}^{l}\in\mathbb{R}^{N\times 1}$. The state size is represented by $N$. The discretization method for computing $\hat{A}$ and $\hat{B}$ is indicated by “discretize”. The vector $h_{t-1}^{{}^{\prime}l}\in\mathbb{R}^{H^{\prime}\times 1}$ and the SSM state is represented by $S_{t}^{l}\in\mathbb{R}^{H^{\prime}\times N}$. The symbol $\otimes$ denotes element-wise multiplication, and $\hat{B}_{t-1}^{l}\left({h^{\prime l}_{t-1}}\right)^{T}$ represents matrix multiplication. It is important to note that the definitions of $A^{l}$ and $\hat{A}^{l}$ presented here differ from their original definitions due to Mamba’s simplification to diagonal matrices.
Mamba improves sequence modelling ability of the state space model by enabling the model to focus on or filter out inputs into a sequential state.
Mamba2
Dao and Gu (2024) theoretically proves the connections between structured state space models and attention mechanisms. They also simplify structured matrix $\hat{A}$ further into scalar-times-identity structure and thus develop a new state space duality (SSD) framework with multi-head patterns similar to transformers. This modification trades in the expressiveness of the $\hat{A}$ matrix for faster training and enlarged state size.
@@
Our proposed method draws some spirits from techniques employed in KV cache compression (Mu, Li, and Goodman 2023; Ge et al. 2024; Yang et al. 2024a; Chevalier et al. 2023; Hwang et al. 2024; Gao, Cao, and Li 2024) by leveraging the language model itself to aggregate information via hidden states and employing a scoring mechanism to select the most salient representations. Nevertheless, different from transformers, ReMamba’s compression strategy focuses on two key objectives: 1) compressing and selectively retaining crucial information to minimize information degradation, and 2) reducing the frequency of state space updates to further alleviate the information loss.
Stage1 : Selective Compression
@@
In Figure 2, the compression hyperparameter settings are: $s=0.2$, $p=0.4$, $range=(0.2,0.6)$, $\mathcal{R}=[3,6]$, $\rho=0.5$, $K=2$. In our experiments, we find that $s=0$ yields the best results, which can be attributed to the casual language modeling nature of Mamba (this will be discussed in more details later).
@@
Stage 2: Selective Adaption
@@
Training
Following the forward encoding processes, standard causal language generation is applied using the Mamba architecture. During training, newly introduced parameters within the selective compression mechanism are optimized. These parameters, except for $\Theta$ which is initialized to all zeros, are initialized with a subset of the weights from the first layer’s in_proj matrix. Additionally, for parameters in Mamba, the dt_proj matrix is fully trained, while in_proj, out_proj, embeddings, and lm_head are updated using Low-Rank Adaptation (LoRA) (Hu et al. 2022). In our best implementation, to emphasize the significance of specific information, gradients flowing into the importance scores are scaled proportionally to these scores. This approach intuitively prioritizes the training of more critical representations.
Model
2WikiMQA
GovReport
HotpotQA
LCC
MultiNews
MultiQA
PassCount
PassRetrie.
Qasper
RepoBench
SAMSum
TREC
TriviaQA
Average
llama2-3b (Pre) 13.24 25.98 12.64 60.20 16.36 27.75 1.03 6.33 8.41 46.92 32.68 52.33 62.43 28.18
llama2-3b (SFT) 17.10 23.79 22.51 57.79 20.27 33.69 0.00 6.67 20.56 43.47 34.62 52.33 49.64 29.42
Mamba (Pre) 3.73 8.72 4.03 24.03 11.31 4.95 0.80 1.75 3.67 12.83 6.86 9.00 17.40 8.39
Mamba (SFT) 22.10 19.08 15.90 40.20 19.36 30.28 0.00 4.67 19.04 36.02 28.30 39.33 45.97 24.63
ReMamba (SFT) 21.18 19.67 20.56 48.21 18.86 26.39 3.21 6.83 16.76 40.40 33.65 48.67 57.73 27.86
Table 1: Performance on LongBench-E (English branch). “MultiQA” denotes MultiFieldQA , “PassCount” denotes PassageCount, “PassRetrie.” denotes PassageRetrieval. Mamba and ReMamba models are evaluated using a maximum length of 6K tokens, matching their training configurations. Llama2-3B is capped at 2K tokens due to positional encoding limitations, which proves to be the best setting. Here “(Pre)” means pretrained model. “(SFT)” means finetuned model.
@@
llama2-3b (Pre) 2k 0.00 19.33 3.00 23.27 56.25 18.22 20.01
llama2-3b (SFT) 2k 2.22 25.00 4.00 27.23 58.59 20.45 22.92
@@
Table 2: Model performance on closed-ended tasks of L-Eval. “Tokens” denotes the max length. “Finetuned” denotes whether the model is finetuned or not. Mamba and ReMamba models are evaluated using a maximum length of 6K tokens, matching their training configuration. In contrast, llama2-3B is capped at 2K tokens due to positional encoding limitations.
Experiments
Experimental Setups
Our model is designed for long-context question-answering tasks, necessitating a substantial corpus of long-context instruction tuning data. To this end, we leverage the OpenOrca dataset (Mukherjee et al. 2023) and LongAlpaca-12k (Chen et al. 2024). The former comprises a rich collection of ChatGPT-augmented FLAN data alignments, while the latter is a long-context alignment dataset. We initially filter long instruction tuning instances from OpenOrca and concatenate them with LongAlpaca. To accommodate device memory constraints, prompts are truncated to a maximum length of 6,000 tokens. This process yields approximately 200,000 long-context training examples. To augment training data diversity, the initial 300,000 standard instances from OpenOrca are incorporated. During training, the hyperparameter $s$ is fixed at 0. The hyperparameter $p$ is randomly sampled from the interval [0.1, 0.3], while $\rho$ is randomly sampled from the interval [0.05, 0.2].
We finetune the baseline Mamaba 2.8b model and our ReMamba model on the same dataset using LoRA with the same hyperparameter setting (r=32, others default). We also finetune a llama2-3b (Geng and Liu 2023) models for reference. Given the 2k maximum positional encoding limit of llama2-3b, we conduct fine-tuning experiments using two configurations: one matching the Mamba data and another with data truncated to a maximum length of 2K. Comparative results reveal superior performance for the truncated model, which is subsequently adopted as the default setting.
Evaluations
We conduct comparative analyses of our model against baseline Mamba2.8b (both of finetuned and pretrained) on the widely adopted LongBench benchmark (Bai et al. 2024) and LEval benchmark (An et al. 2023), which encompass a diverse set of challenging real-world long-context tasks. For consistency, the same prompt templates and greedy decoding configurations are employed across all models. To provide a reference point, the performance of a similarly sized transformer architecture (llama2-3b) is also included. It is important to note that due to llama2-3b’s 2k maximum positional encoding limitation, its performance in the 6k setting is subpar, necessitating the utilization of its optimal 2k configuration for evaluation.
Results
Model
2WikiMQA
GovReport
HotpotQA
LCC
MultiNews
MultiQA
PassCount
PassRetrie.
Qasper
RepoBench
SAMSum
TREC
TriviaQA
Average
Mamba2 (Pre) 2.18 4.76 1.54 23.46 7.71 2.88 0.60 1.47 1.17 14.97 2.07 8.33 10.60 6.29
Mamba2 (SFT) 13.73 19.97 15.05 41.78 19.52 25.20 0.67 5.67 13.44 38.79 33.95 49.67 43.54 24.69
ReMamba2 (SFT) 18.90 19.03 18.09 51.15 17.68 25.82 3.33 5.00 14.84 43.99 23.55 44.00 56.90 26.33

### v4

@@ Abstract
ReMamba: Equip Mamba with Effective Long-Sequence Modeling
Danlong Yuan1,2thanks: Work done during internship at Meituan, Jiahao Liu5, Bei Li5, Huishuai Zhang1,3thanks: Corresponding author., Jingang Wang5, Xunliang Cai5 Dongyan Zhao1,2,3,422footnotemark: 2 1Wangxuan Institute of Computer Technology, Peking University, 2Center for Data Science, AAIS, Peking University, 3National Key Laboratory of General Artificial Intelligence, 4BIGAI, Beijing, China;, 5Meituan Correspondence: danlongyuan@stu.pku.edu.cn, {zhanghuishuai,zhaodongyan}@pku.edu.cn, {liujiahao12,libei17,wangjingang02,caixunliang}@meituan.com Code: https://github.com/lblankl/ReMamba
1 Introduction
@@
Transformers Vaswani et al. (2017), which form the backbone of most LLMs, encounter substantial challenges when dealing with long texts. The quadratic computational demands and the linear memory costs of the attention mechanism become prohibitive as the text length grows. This complexity poses a significant barrier to effectively modeling long texts, which is crucial for the development of LLMs. To address this, Mamba is proposed as a solution Gu and Dao (2024), which utilizes a recurrent inference mode that ensures linear time complexity and compress information into the fixed state size. This results in constant memory demands during inference. Furthermore, Mamba eliminates the need for positional encoding, theoretically allowing it to handle inputs of any length, while performing competitively against transformers on downstream tasks. Shortly after, Mamba2 was introduced, simplifying the structured $A$ matrix of Mamba to enable faster training and enlarged state size Dao and Gu (2024).
@@
Figure 1: A comparison of pretrained Mamba models and Transformers of equivalent size across speed, short-context, and long-context performance metrics. Speed is measured under conditions of 6k input tokens and 1k output tokens. “short scores” represents the average accuracy across six tasks (HellaSwag, PIQA, Arc-E, Arc-C, WinoGrande, OpenbookQA) evaluated within the LM evaluation harness Gao et al. (2023). “long scores” corresponds to the average scores on the LongBench-E benchmark Bai et al. (2024). Notably, all LongBench evaluations employ a maximum token length of 2k to align with the model’s training configuration.
@@
Despite these advantages, some studies reveal that Mamba models do not perform as well as expected when dealing with long texts reaching 2k tokens or more Waleffe et al. (2024). As depicted in Figure 1, our experimental findings reveal that the pretrained Mamba model surpasses pretrained Transformers of comparable size, such as Llama-3b  Geng and Liu (2023), on short-context tasks. Conversely, a substantial performance degradation is observed for Mamba on long-context tasks relative to Transformers. This performance disparity underscores a significant limitation of Mamba models in practical long-context applications.
@@
This long-context deficency issue of Mamba is usually attributed to its RNN-like nature. This kind of architecture exhibits limitations in preserving crucial information from earlier input sequences as the context length increases due to the fixed-size memory Wen et al. (2024); Yang et al. (2024b). Hybrid architectures Lieber et al. (2024); Ren et al. (2024); Park et al. (2024) have sought to mitigate this issue by integrating attention mechanisms from transformers. However, these approaches often lead to decreased computational efficiency and increased memory consumption.
@@
The key challenge in Mamba is the excessive degradation of distant information. ReMamba addresses this by employing an effective compression strategy that condenses the information and reduces the context length. Specifically, it selects the top-k hidden states during the first forward pass and integrates them into the state space using Mamba’s selective mechanism in the second pass. ReMamba introduces minimal computational overhead (one additional forward pass) and maintains low, constant memory consumption. Experimental results demonstrate that our approach significantly improves Mamba’s long-context performance, bringing it close to the performance of transformers. Our ReMamba model achieves a 3.2 improvement over the baseline on LongBench Bai et al. (2024) and 1.6 improvement on L-Eval An et al. (2023). Furthermore, our methodology exhibits transferability to Mamba2, yielding a 1.6 improvement on LongBench.
2 Related work
2.1 Mamba
@@
The state space model chooses the time-invariant $\hat{A}$ (state transition matrix) and $\hat{B}$ (input coefficient matrix) thus lacking expressiveness and flexibility. Mamba Gu and Dao (2024) proposes to make $\hat{A}$ and $\hat{B}$ dynamically depend on inputs.
@@
Furthermore, $\Delta_{t-1}^{l}\in\mathbb{R}^{H^{\prime}}$ is the discrete time step corresponding to the selective mechanism in Mamba, where $H^{\prime}$ is the intermediate hidden size. The continuous and discrete state transformation matrices at layer $l$ are given by $A^{l},\hat{A}^{l}\in\mathbb{R}^{H^{\prime}\times N}$, respectively. The continuous and discrete input coefficient matrices are denoted by $B_{t-1}^{l},\hat{B}_{t-1}^{l}\in\mathbb{R}^{N\times 1}$. The state size is represented by $N$. The discretization method for computing $\hat{A}$ and $\hat{B}$ is indicated by “discretize”. The vector $h_{t-1}^{{}^{\prime}l}\in\mathbb{R}^{H^{\prime}\times 1}$ and the SSM state is represented by $S_{t}^{l}\in\mathbb{R}^{H^{\prime}\times N}$. The symbol $\otimes$ denotes element-wise multiplication, and $\hat{B}_{t-1}^{l}\left({h^{\prime l}_{t-1}}\right)^{T}$ represents matrix multiplication.
2.2 Mamba2
Dao and Gu (2024) theoretically proves the connections between structured state space models and attention mechanisms. They also simplify structured matrix $\hat{A}$ further into scalar-times-identity structure and thus develop a new state space duality (SSD) framework with multi-head patterns similar to transformers.
2.3 Long Context Mamba and Transformers
Positional interpolation has been widely used as a technique to extend the context length of transformers Chen et al. (2023); Peng et al. (2024); Ding et al. (2024). But they are specialized for transformers.
Mamba has been found to struggle in maintaining performance beyond its pretraining context length without additional training. LongMamba Peiyuan (2024) made the first successful attempt to extend Mamba’s context length through a few hours of long-context fine-tuning. DeciMamba Ben-Kish et al. (2024) aimed to address the context extension problem of Mamba in a training-free manner, proposing a method to progressively reduce sequence length across layers by empirically removing unimportant tokens.
However, our experiments demonstrate that long-context fine-tuned Mamba still lags behind long-context fine-tuned transformers of the same size, despite using the same data. Moreover, DeciMamba2.8b appears to be insufficiently effective when evaluated on two widely used long-context benchmarks.
3 Preliminary Study
KV cache compression is widely used in transformers to reduce memory consumption and improve inference speed. However, prompt compression often results in performance degradation compared to the full context lengths generation in transformers. Unlike transformers, Mamba does not employ a KV cache; instead, it utilizes a fixed-size state space in each layer to preserve context memory. A potential issue with Mamba is its tendency to forget distant information. In our preliminary study, we hypothesize that the state space update in Mamba is insufficient for effectively compressing context information, and that techniques like prompt compression could help relieve this issue.
To explore this, we apply a simple prompt compression method: we replace part of the context tokens with a few randomly selected hidden states from the last layer of Mamba, creating a shorter prompt (referred to as random Mamba). This approach is similar to that of Ge et al. (2024), which used soft prompts for information compression. Intuitively, this random compression method may lead to significant information loss and degraded performance. However, our results show that the average scores for different context lengths on LongBench between normal Mamba and random Mamba are similar when both are trained on the same long-context dataset. Furthermore, random Mamba outperforms normal Mamba at certain context lengths, as shown in Figure 3 of 5.4.1. The random_select (SFT) represents the fine-tuned random Mamba, while Mamba (SFT) represents the fine-tuned vanilla Mamba.
This observation suggests that information loss in Mamba when handling long contexts is substantial. To relieve this, we propose selective compression and selective adaptation through leveraging Mamba’s state space update mechanism.
@@ 4 Methodology
Our proposed method draws some spirits from techniques employed in KV cache compression Mu et al. (2023); Ge et al. (2024); Yang et al. (2024a); Chevalier et al. (2023); Hwang et al. (2024); Gao et al. (2024) by leveraging the language model itself to aggregate information via hidden states and employing a scoring mechanism to select the most salient representations. Nevertheless, different from transformers, ReMamba’s compression strategy focuses on two key objectives: 1) compressing and selectively retaining crucial information to minimize information degradation, and 2) reducing the frequency of state space updates to further alleviate the information loss.
The selection method employed in ReMamba is a simplified key-query-value-based approach, commonly used in Retrieval-Augmented Generation and summarization tasks Lewis et al. (2020); Mao et al. (2022). In this context, we select the most important hidden states from the last layer of Mamba to mitigate information loss during the compression process.
4.1 Stage1 : Selective Compression
@@
In Figure 2, the compression hyperparameter settings are: $s=0.2$, $p=0.4$, $range=(0.2,0.6)$, $\mathcal{R}=[3,6]$, $\rho=0.5$, $K=2$. In our experiments, we find that $s=0$ yields the best results, which can be attributed to the casual language modeling nature of Mamba (this will be discussed in more details in Appendix A.3).
@@
4.2 Stage 2: Selective Adaption
@@
4.3 Training
Following the forward encoding processes, standard causal language generation is applied using the Mamba architecture. During training, newly introduced parameters within the selective compression mechanism are optimized. These parameters, except for $\Theta$ which is initialized to all zeros, are initialized with a subset of the weights from the first layer’s in_proj matrix. Additionally, for parameters in Mamba, the dt_proj matrix is fully trained, while in_proj, out_proj, embeddings, and lm_head are updated using Low-Rank Adaptation (LoRA) Hu et al. (2022). In our best implementation, to emphasize the significance of specific information, gradients flowing into the importance scores are scaled proportionally to these scores. This approach intuitively prioritizes the training of more critical representations.
Model 2WikiMQA GovReport HotpotQA LCC MultiNews MultiQA PassCount PassRetrie. Qasper RepoBench SAMSum TREC TriviaQA Average Llama-3b (Pre) 4.07 10.33 3.93 43.17 6.96 7.74 1.78 7.13 3.49 32.14 9.98 15.0 30.65 13.57 Llama-3b (SFT) 15.69 24.55 19.69 56.17 19.37 29.73 0.33 6.67 19.73 44.47 33.37 48.67 58.41 28.99 DeciMamba (Pre) 3.89 9.36 3.85 25.69 11.62 4.91 0.83 1.19 3.21 13.88 6.79 6.67 17.72 8.43 DeciMamba (SFT) 19.6 13.74 15.32 23.65 10.91 17.88 1.37 4.96 5.72 12.54 10.59 30.33 46.06 16.36 Mamba (Pre) 3.73 8.72 4.03 24.03 11.31 4.95 0.80 1.75 3.67 12.83 6.86 9.00 17.40 8.39 Mamba (SFT) 22.10 19.08 15.90 40.20 19.36 30.28 0.00 4.67 19.04 36.02 28.30 39.33 45.97 24.63 ReMamba (SFT) 21.18 19.67 20.56 48.21 18.86 26.39 3.21 6.83 16.76 40.40 33.65 48.67 57.73 27.86
Table 1: Performance on LongBench-E (English branch). “MultiQA” denotes MultiFieldQA , “PassCount” denotes PassageCount, “PassRetrie.” denotes PassageRetrieval. Models are evaluated using a maximum length of 6K tokens, matching their finetuning configurations. Here “(Pre)” means pretrained model. “(SFT)” means finetuned model.
@@ 4.3 Training
Llama-3b (Pre) 6k 0.0 24.71 3.0 27.23 32.81 17.47 17.54
Llama-3b (SFT) 6k 1.11 19.62 7.0 24.26 57.03 27.14 22.69
DeciMamba (Pre) 6k 2.22 24.71 0.0 25.25 24.6 16.73 15.59
DeciMamba (SFT) 6k 0.0 23.69 0.0 26.86 53.17 22.68 21.07
@@
Table 2: Model performance on closed-ended tasks of L-Eval. “Tokens” denotes the max length. “SFT” denotes finetuned models. “Pre“ denotes pretrained models.
5 Experiments
5.1 Experimental Setups
Our model is designed for long-context question-answering tasks, necessitating a substantial corpus of long-context instruction tuning data. To this end, we leverage the OpenOrca dataset Mukherjee et al. (2023) and LongAlpaca-12k Chen et al. (2024). The former comprises a rich collection of ChatGPT-augmented FLAN data alignments, while the latter is a long-context alignment dataset. We initially filter long instruction tuning instances from OpenOrca and concatenate them with LongAlpaca. To accommodate device memory constraints, prompts are truncated to a maximum length of 6,000 tokens. This process yields approximately 200,000 long-context training examples. To augment training data diversity, the initial 300,000 standard instances from OpenOrca are incorporated. This dataset is referred to as the LongOrca dataset. We finetune the baseline Mamaba 2.8b model and our ReMamba model on the same dataset. We also finetune a DeciMamba2.8b Ben-Kish et al. (2024) and a Llama-3b Geng and Liu (2023) for reference. DeciMamba aims to address Mamba’s context-extension issue without requiring additional training. Although our approach differs slightly in terms of settings and objectives, we still fine-tune DeciMamba2.8b using the same data. Given the 2k maximum positional encoding limit of Llama-3b, we conduct fine-tuning experiments using the simple linear positional interpolation technique Chen et al. (2023) to extend its context length. The data construction process for Llama-3b is identical to that of Mamba. Details can be found in A.1. Notice that Mamba2.8b is the largest model we can obtain.
5.2 Evaluations
We conduct comparative analyses of our model against baseline Mamba2.8b and DeciMamba2.8b (both of finetuned and pretrained) on the widely adopted LongBench benchmark Bai et al. (2024) and LEval benchmark An et al. (2023), which encompass a diverse set of challenging real-world long-context tasks. For consistency, the same prompt templates and greedy decoding configurations are employed across all models.
To provide a reference point, the performance of a similarly sized transformer architecture (Llama-3b) is also included. Both the pretrained and fine-tuned evaluations of Llama-3b utilize the linear positional interpolation technique.
5.3 Results